OpenStudy (anonymous):

(sqrt)9y4 + 12 (sqrt) y6 - 3 (sqrt)y6 (sqrt) y4

jhonyy9 (jhonyy9):

3y^2 +12y^3 -3y^3 what sign you have missed there ??? y^2

jhonyy9 (jhonyy9):

or there is multiplication ?

OpenStudy (anonymous):

\[\sqrt{9y ^{4}} + 12\sqrt{y ^{6}} - 3\sqrt{y ^{6}} -\sqrt{y ^{4}}\]

jhonyy9 (jhonyy9):

ok than you need to put the sign minus 3y^3 ok ?

OpenStudy (anonymous):

huh?